Heaton Steel Company is a steel manufacturer. It has three departments. Department A is responsible for melting the scrap metal. Afterwards, the molten metals are transferred to Department B. Department B will then pours the molten metal into ingot casting. Ingot castings are transferred to Department C. Department C sands off any sharp edges and package the castings for shipment to customers. Data for Department A for January are:

Department A: Costs of Beginning Work-in-Process (January 1st)

  • Scrap Metal (5,000 tons)                                            $6,000
  • Labor work for melting the metal                              $5,000
  • Factory Overhead                                                       $500

Department A: Costs Added in January

  • Scrap Metal (67,000 tons)                                          $80,000
  • Labor                                                                          $130,000
  • Factory Overhead                                                       $4,500

Additional information

  • 68,000 tons of molten metals transferred to Department B
  • January 31st: 50% complete as to conversion costs

Complete the Cost Production Report Below

ç Inputs ➔

=

ç Outputs ➔

Unit Calculations

Beginnin g WIP Inventory

Add Current Month

Equals Total Production

Completed Transfer Out

Add Ending WIP Inventory

Total Outputs

Direct Materials Whole Units

% Equivalent Completed

Conversion Equivalent Units

ç Inputs ➔

Cost of Production per Unit

ç Outputs ➔

Cost Calculations

Beginnin g WIP Inventory

Add Current Month

Equals Total Production

Completed Transfer Out

Add Ending WIP Inventory

Total Outputs

Direct Materials Units

Direct Materials Cost

Conversion Units

Direct Labor

Manufacturing Overhead

Total Conversion Cost

Total
Answer #1

Solutions:

Complete the Cost Production Report Below
ç Inputs ➔ = ç Outputs ➔
Unit Calculations Beginnin g WIP Inventory Add Current Month Equals Total Production Completed Transfer Out Add Ending WIP Inventory Total Outputs
Direct Materials Whole Units 5000 67000 72000 68000 4000 72000
% Equivalent Completed 100% 50%
Conversion Equivalent Units 68000 2000 70000
ç Inputs ➔ Cost of Production per Unit ç Outputs ➔
Cost Calculations Beginnin g WIP Inventory Add Current Month Equals Total Production Completed Transfer Out Add Ending WIP Inventory Total Outputs
Direct Materials Units 5000 67000 72000 68000 4000 72000
Direct Materials Cost 6000 80000 86000 1.2 81222 4778 86000
Conversion Units 68000 2000
Direct Labor 5000 130000 135000 1.93 131143 3857 135000
Manufacturing Overhead 500 4500 5000 0.07 4857 143 5000
Total Conversion Cost 5500 134500 140000 2.00 136000 4000 140000
Total 11500 214500 226000 217222 8778 226000

Notes:

Material cost per unit is calculated as Total Costs / Total units = 86000/72000

Direct Labor and Overhead costs = Total Costs / Equivalent Units = 135000/70000 and 5000/70000

Equivalent Units = 50% complete as to conversion costs = 68000 + 50%*4000 = 70000